b) An escalator at an airport has an acceleration function

a       =      [6.33t2 i  - 16.8t3 j] mm s-2

         Determine the escalator’s velocity 1.2 s after a person steps on it.  

                  

c)    If a force vector F = {12i + 13j – 14k}N, what is its unit vector?

 

Q4    a)     Distinguish between systematic errors and random errors.

b)     What is s and its uncertainty, if s = a3bc1/2 / d, where:

         a       =      (3.00  0.150) m

         b      =      (2.30  0.30) s2

         c       =      (49  0.3) kg4

         d      =      (150  10.5) m

1
Expert's answer
2021-12-20T10:28:23-0500
"\vec{v}=6.33\frac{t^3}{3}\vec{i}-16.8\frac{t^4}{4}\vec{j}\\
\vec{v}(2)=6.33\frac{2^3}{3}\vec{i}-16.8\frac{2^4}{4}\vec{j}\\\vec{v}(2)=(16.9\vec{i}-67.2\vec{j})\frac{mm}{s}"